OpenStudy (anonymous):

add simplify if possible: 7/y-1 + 4/(y-1)^2 =

OpenStudy (anonymous):

what is a common denominator equal to ?

OpenStudy (anonymous):

not understanding your question

OpenStudy (anonymous):

do you know the definition of numerator and denominator? what is the common denominator means?

OpenStudy (anonymous):

the denominators are y-1 and y-2^2

OpenStudy (anonymous):

so, to be able to add two fractions you need COMMON denominator, right? what is the common denominator in this case?

OpenStudy (anonymous):

correction y-1 and y-1^2

OpenStudy (anonymous):

i got that... happens sometime

OpenStudy (anonymous):

y=-2

OpenStudy (anonymous):

so, one fraction has denominator: (y-1) another: (y-1)^2, which is = (y-1)*(y-1) right?

OpenStudy (anonymous):

break them down into 2 equeations: 7/y-1 abd 4(y-1)^2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok then what this stuff is new to me

OpenStudy (anonymous):

common demonotator for 7/y-1 is y. since y x 1 is 1y they cancel out so its just 7

OpenStudy (anonymous):

=\[=[7(y-1) +4]/(y-1)^{2}=(7y-3)/(y-1)^{2}\]

OpenStudy (anonymous):

when you have a number before the bracketsd it means you have to multiply every number by that number so it would be (4y-4)^2

OpenStudy (anonymous):

so its (4y-4)(4y-4) and you should know it from there

OpenStudy (anonymous):

biljana92000 ... what are you talking about...? unless I'm reading the original problem wrong. jazi1, please check the problem & let me know if this is right: \[7/(y-1) +4/(y-1)^{2}=...\] did i read it wrong?

OpenStudy (anonymous):

that's correct inik but the first y-1 is not in parenthese

OpenStudy (anonymous):

do you mean its 7 divided by y after that -1 or 7 divided by (y-1)?

OpenStudy (anonymous):

it just like you have but the first (y-1) is not in parenthese but the second one (y-1)^2 is

OpenStudy (anonymous):

again, parentheses or not - different people on web using shortcuts in writing problems & it's OK . I just trying to understand which problem you are trying to solve... it will have VERY different results 1) \[7/y +4/(y-1)^2 -1\] (I change the order just to see better the problem) 2)\[7/(y-1) +4/(y-1)^2\] which means that 7 is divided by y-1, not by1which one is your problem - 1st or 2nd?

OpenStudy (anonymous):

if second - you already have the answer above if first... I'll post it

OpenStudy (anonymous):

7/(y-1) + 4/(y-1)^2

OpenStudy (anonymous):

so inik the answer is -2